Question 7
A streaming TV service charges $15 per month. In terms of monthly costs to serve a customer, they spend $2 in advertising, $1 in variable cost and $1 in fixed cost. The average customer will maintain the service for a year and a half. What is the CLV of their customers assuming a 12% discount rate and 60% retention?

option

A 247.5

b 23.76

c 174.26

d 227.7
Answer #1

CALCULATION OF CUSTOMER LIFETIME VALUE (CLV)

Rate of discount=12%=0.12

Retention rate= 60%=0.6

Average Profit Margin=(15-2-1-1)/15=11/15=0.73333

Average Gross Margin Per Customer Lifespan=$15*18*0.73333=$198

CLV=Average Gross Margin Per Customer Lifespan X (Retention / (1 + Rate Of Discount – Retention))

CLV=$198*(0.6/(1+0.12-0.6))

CLV=$198*1.15

CLV=$227.7

ANSWER:

d. 227.7